# Chirped Gaussians Have Maximal Frame Sets

### v1

@@ 1 Introduction
A Gabor frame on $L^{2}(\mathbb{R})$ is a Gabor system that forms a frame on $L^{2}(\mathbb{R})$. The set of parameters $(\alpha,\beta)$ that makes $\mathcal{G}(g,\alpha\mathbb{Z}\times\beta\mathbb{Z})$ a frame is called the frame set of $g$ and denoted by $F(g)$. The problem of determining $F(g)$ for an arbitrary function $g$ is a long standing difficult open problem in this field. Even for simple window functions such as $\chi_{[0,1)}$, the corresponding frame set could be fairly complicated (see [13]). The classical density theory points out that if $\alpha\beta>1$, then $\mathcal{G}(g,\alpha\mathbb{Z}\times\beta\mathbb{Z})$ is incomplete and thus can not possibly be a frame (see e.g., [9, Theorem 3a] or [1]).
The critical case $\alpha\beta=1$ is in general not so interesting since such windows are already completely characterized by the Zak transform. More importantly it has been excluded from frame sets of functions with good time-frequency concentration. This is due to the fact that in such case $\mathcal{G}(g,\alpha\mathbb{Z}\times\beta\mathbb{Z})$ is a frame if and only if the Zak transform of $g$ is essentially bounded away from both 0 and $\infty$ (see e.g., [5, Corollary 8.3.2]), and it is well known that if the transformed function is continuous, then it always admits zeros (see e.g., [5, Lemma 8.4.2]) and thus can not be lowered bounded. A class of functions that will guarantee continuity after the Zak transform is the set of continuous Wiener functions. Wiener functions on $\mathbb{R}$ form a Banach space equipped with the norm (see e.g., [5, Chapter 6.1]).
@@
We shall say that a function $g$ has maximal frame set (or $F(g)$ is maximal), if $F(g)$ contains all pairs $(\alpha,\beta)$ with $\alpha\beta<1$ and $\alpha,\beta>0$. A primal question is then to identify all $W_{0}(\mathbb{R})$ functions whose frame sets are maximal. It was first conjectured in [3] that Gaussians possess such a property. This is proved in [19, 23, 24]. Later the hyperbolic secant $1/(e^{x}+e^{-x})$ [14], the one sided exponential $e^{-x}\chi_{\mathbb{R}^{+}}(x)$ [11] (actually this is not a $W_{0}(\mathbb{R})$ function since it is not even continuous, its Zak transform has no zero and its frame set contains also the critical case $\alpha\beta=1$, see [13, Section 3.2.1]) and the two sided exponential $e^{-|x|}$ [12] are also found to have this property. Then apparently automorphisms of the $L^{2}$ space such as translations, modulations, dilations, scalings and Fourier transforms of these functions will inherit this property.
@@
where $c>0$, $\gamma\geq 0$, $\nu,\nu_{j}\in\mathbb{R}$, $N\in\mathbb{N}\cup\{\infty\}$ and $0<\gamma+\sum_{j=1}^{N}\nu_{j}^{2}<\infty$. This is characterized in [20, 21, 22]. In a series of papers [6, 7, 8] it is proved that if $N$ is finite or if $N=\infty$ but $\alpha\beta\in\mathbb{Q}$, then $\alpha\beta<1$ is in the frame set of $g$ (in particular, if $g$ is at the same time in the Feichtinger algebra, then $F(g)$ is open [4], and thereof maximal), which covers and also greatly extends all aforementioned known cases. So far these are the only known functions to have maximal frame sets. Mysteriously the Zak transforms of all these functions except the one-sided exponential possess a unique simple zero in the unit square (see [2, 16, 17, 25]).
Is total positivity necessary for maximality of frame sets? If we modulate a totally positive function, or simply multiply it by a negative number, then it is not totally positive any more but its frame set stays invariant. Therefore to discuss necessity it perhaps makes sense to phrase it as “is total positivity necessary for real non-negative Wiener functions to have maximal frame sets?”
The purpose of this paper is to produce complex valued $W_{0}(\mathbb{R})$ functions whose frame sets are maximal and whose Zak transforms also have unique simple zeros at center of the unit square. It is clear from their expressions that these functions are neither totally positive nor time-frequency shifts or dilations or Fourier transforms of totally positive functions. Let
@@ 3.1 The fractional Fourier transform and chirps
These properties are well known, see perhaps [18, Section 4] and references therein for a concise exposition of properties of $\mathcal{F}^{(\theta)}$ tailored to time-frequency analysis.
@@
where the term $e^{-\pi i/4}$ comes from taking the reciprocal of a square root of $i$, while if $\lambda^{\prime}<0$, then we have to take it on i-i and the formula becomes (alternatively this can also be obtained by applying $\mathcal{F}$ again, see also the table in [15, A-7])
@@ 3.3 Maximality: the general case
Recall the formula (see [15, A-7]) that if $\alpha,\beta\in\mathbb{R}$, $\alpha+\beta i\neq 0$ and $\alpha\geq|\beta|$, then
@@
For the necessity, that $0<|\det Q|\leq 1$ is necessary for $\mathcal{G}(\varphi,Q\mathbb{Z}^{2})$ to be a frame is asserted in [9, Theorem 10]. For the critical case of $1=|\det Q|=|\alpha\beta|$ we see that the above derivation actually indicates that $\mathcal{G}(\varphi,Q\mathbb{Z}^{2})$ is a frame if and only if $\mathcal{G}(h_{\lambda}\cdot\varphi,D_{\alpha,\beta}\mathbb{Z}^{2})$ is so. But $h_{\lambda}\cdot\varphi\in W_{0}(\mathbb{R})$. Thus by the Amalgam Balian-Low theorem it can not be a frame if $|\alpha\beta|=1$. ∎
@@ 4.1 The Zak transform and the theta function
which indicates that the values of $Zf$ on $\mathbb{R}^{2}$ is completely determined by its values on $S$. See [5, Chapter 8] for more of its properties.
@@
This form of the theta function is closely related to the Zak transform of Gaussians (see e.g.,[14]). In this article the following decomposition is relevant ([2, Lemma 2.1], which is also available in [26, Chapter 21.3, p.469]):
@@
then we have ([26, Chapter 21.3, p.469])
@@ 4.2 Location, uniqueness, and simplicity of zeros
The Zak transform method is particularly effective in analyzing frame sets of the critical case and other integer oversampling cases since in such cases it diagonalizes the frame operator and makes it a multiplication operator [5, Corollary 8.3.2, Equation 8.21], i.e., if $\alpha=1/n$ with $n\in\mathbb{N}$, then $\mathcal{G}(g_{\gamma},\alpha\mathbb{Z}\times\mathbb{Z})$ is a frame if and only if
@@
The other cases are more complicated. But in general they can be characterized as ([10], also stated as [7, Theorem 2.3]): $\mathcal{G}(g_{\gamma},\alpha\mathbb{Z}\times\mathbb{Z})$ is a frame if and only if
@@
(a) is a classical result, and we refer to [5, Lemma 8.4.2] for a proof and references therein for its origin. (f) is [13, Section 3.2.1].
@@
For (e), suppose that $\lambda\neq-i$ is the eigenvalue. Then we simply apply the formula ([5, Proposition 8.2.2]) that
@@
(1) and (2) are in [13, Section 3.2.2-3.2.3]. (3) is established in [2] for Gaussians, in [16] for exponential $B$-splines of finite order and for finite type totally positive functions without the Gaussian factor, in [17] for infinite type totally positive functions without the Gaussian factor, and finally in [25] for totally positive functions with the Gaussian factor and for their convolutions with exponential $B$-splines.

### v2

@@ 1 Introduction
A Gabor frame on $L^{2}(\mathbb{R})$ is a Gabor system that forms a frame on $L^{2}(\mathbb{R})$. The set of parameters $(\alpha,\beta)$ that makes $\mathcal{G}(g,\alpha\mathbb{Z}\times\beta\mathbb{Z})$ a frame is called the frame set of $g$ and denoted by $F(g)$. The problem of determining $F(g)$ for an arbitrary function $g$ is a long standing difficult open problem in this field. Even for simple window functions such as $\chi_{[0,1)}$, the corresponding frame set could be fairly complicated (see [14]). The classical density theory points out that if $\alpha\beta>1$, then $\mathcal{G}(g,\alpha\mathbb{Z}\times\beta\mathbb{Z})$ is incomplete and thus can not possibly be a frame (see e.g., [10, Theorem 3a] or [1]).
The critical case $\alpha\beta=1$ is in general not so interesting since such windows are already completely characterized by the Zak transform. More importantly it has been excluded from frame sets of functions with good time-frequency concentration. This is due to the fact that in such case $\mathcal{G}(g,\alpha\mathbb{Z}\times\beta\mathbb{Z})$ is a frame if and only if the Zak transform of $g$ is essentially bounded away from both 0 and $\infty$ (see e.g., [6, Corollary 8.3.2]), and it is well known that if the transformed function is continuous, then it always admits zeros (see e.g., [6, Lemma 8.4.2]) and thus can not be lowered bounded. A class of functions that will guarantee continuity after the Zak transform is the set of continuous Wiener functions. Wiener functions on $\mathbb{R}$ form a Banach space equipped with the norm (see e.g., [6, Chapter 6.1]).
@@
We shall say that a function $g$ has maximal frame set (or $F(g)$ is maximal), if $F(g)$ contains all pairs $(\alpha,\beta)$ with $\alpha\beta<1$ and $\alpha,\beta>0$. A primal question is then to identify all $W_{0}(\mathbb{R})$ functions whose frame sets are maximal. It was first conjectured in [4] that Gaussians possess such a property. This is proved in [20, 24, 25]. Later the hyperbolic secant $1/(e^{x}+e^{-x})$ [15], the one sided exponential $e^{-x}\chi_{\mathbb{R}^{+}}(x)$ [12] (actually this is not a $W_{0}(\mathbb{R})$ function since it is not even continuous, its Zak transform has no zero and its frame set contains also the critical case $\alpha\beta=1$, see [14, Section 3.2.1]) and the two sided exponential $e^{-|x|}$ [13] are also found to have this property. Then apparently automorphisms of the $L^{2}$ space such as translations, modulations, dilations, scalings and Fourier transforms of these functions will inherit this property.
@@
where $c>0$, $\gamma\geq 0$, $\nu,\nu_{j}\in\mathbb{R}$, $N\in\mathbb{N}\cup\{\infty\}$ and $0<\gamma+\sum_{j=1}^{N}\nu_{j}^{2}<\infty$. This is characterized in [21, 22, 23]. In a series of papers [7, 8, 9] it is proved that if $N$ is finite or if $N=\infty$ but $\alpha\beta\in\mathbb{Q}$, then $\alpha\beta<1$ is in the frame set of $g$ (in particular, if $g$ is at the same time in the Feichtinger algebra, then $F(g)$ is open [5], and thereof maximal), which covers and also greatly extends all aforementioned known cases. Mysteriously the Zak transforms of all these functions except the one-sided exponential possess a unique simple zero in the unit square (see [2, 17, 18, 26]).
Is total positivity necessary for maximality of frame sets? If we modulate a totally positive function, or simply multiply it by a negative number, then it is not totally positive any more but its frame set stays invariant. Therefore to discuss necessity it perhaps makes sense to phrase it as “is total positivity necessary for real non-negative Wiener functions to have maximal frame sets?” In fact it is pointed out in [3] that certain linear combinations of Cauchy kernels do admit maximal frame sets, these functions are neither totally positive nor (compositions of) time-frequency shifts or dilations or Fourier transforms of totally positive functions, but they are not Wiener functions either.
The purpose of this paper is to produce a class of complex valued $W_{0}(\mathbb{R})$ functions whose frame sets are maximal and whose Zak transforms also have unique simple zeros at center of the unit square. It is clear from their expressions that these functions are as well neither totally positive nor (compositions of) time-frequency shifts or dilations or Fourier transforms of totally positive functions. Let
@@ 3.1 The fractional Fourier transform and chirps
These properties are well known, see perhaps [19, Section 4] and references therein for a concise exposition of properties of $\mathcal{F}^{(\theta)}$ tailored to time-frequency analysis.
@@
where the term $e^{-\pi i/4}$ comes from taking the reciprocal of a square root of $i$, while if $\lambda^{\prime}<0$, then we have to take it on i-i and the formula becomes (alternatively this can also be obtained by applying $\mathcal{F}$ again, see also the table in [16, A-7])
@@ 3.3 Maximality: the general case
Recall the formula (see [16, A-7]) that if $\alpha,\beta\in\mathbb{R}$, $\alpha+\beta i\neq 0$ and $\alpha\geq|\beta|$, then
@@
For the necessity, that $0<|\det Q|\leq 1$ is necessary for $\mathcal{G}(\varphi,Q\mathbb{Z}^{2})$ to be a frame is asserted in [10, Theorem 10]. For the critical case of $1=|\det Q|=|\alpha\beta|$ we see that the above derivation actually indicates that $\mathcal{G}(\varphi,Q\mathbb{Z}^{2})$ is a frame if and only if $\mathcal{G}(h_{\lambda}\cdot\varphi,D_{\alpha,\beta}\mathbb{Z}^{2})$ is so. But $h_{\lambda}\cdot\varphi\in W_{0}(\mathbb{R})$. Thus by the Amalgam Balian-Low theorem it can not be a frame if $|\alpha\beta|=1$. ∎
@@ 4.1 The Zak transform and the theta function
which indicates that the values of $Zf$ on $\mathbb{R}^{2}$ is completely determined by its values on $S$. See [6, Chapter 8] for more of its properties.
@@
This form of the theta function is closely related to the Zak transform of Gaussians (see e.g.,[15]). In this article the following decomposition is relevant ([2, Lemma 2.1], which is also available in [27, Chapter 21.3, p.469]):
@@
then we have ([27, Chapter 21.3, p.469])
@@ 4.2 Location, uniqueness, and simplicity of zeros
The Zak transform method is particularly effective in analyzing frame sets of the critical case and other integer oversampling cases since in such cases it diagonalizes the frame operator and makes it a multiplication operator [6, Corollary 8.3.2, Equation 8.21], i.e., if $\alpha=1/n$ with $n\in\mathbb{N}$, then $\mathcal{G}(g_{\gamma},\alpha\mathbb{Z}\times\mathbb{Z})$ is a frame if and only if
@@
The other cases are more complicated. But in general they can be characterized as ([11], also stated as [8, Theorem 2.3]): $\mathcal{G}(g_{\gamma},\alpha\mathbb{Z}\times\mathbb{Z})$ is a frame if and only if
@@
(a) is a classical result, and we refer to [6, Lemma 8.4.2] for a proof and references therein for its origin. (f) is [14, Section 3.2.1].
@@
For (e), suppose that $\lambda\neq-i$ is the eigenvalue. Then we simply apply the formula ([6, Proposition 8.2.2]) that
@@
(1) and (2) are in [14, Section 3.2.2-3.2.3]. (3) is established in [2] for Gaussians, in [17] for exponential $B$-splines of finite order and for finite type totally positive functions without the Gaussian factor, in [18] for infinite type totally positive functions without the Gaussian factor, and finally in [26] for totally positive functions with the Gaussian factor and for their convolutions with exponential $B$-splines.